OpenStudy (anonymous):

If we have that x+2y=20 then the largest possible value for xy is what Maximum product?

OpenStudy (anonymous):

let the product be P = xy. since x+2y=20, solve for one these variables (x or y). then plug this expression into the product equation. what u got now?

OpenStudy (anonymous):

x=20-2y and y=(20-x/2)

OpenStudy (anonymous):

one of 'em will suffice.. let's choose that first one, x=20-2y. so plugging that in to the product equation, P = xy, we get: P = (20-2y)y or P = 20y - 2y^2 understand up to this point?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so you need to maximize this P = 20y-2y^2. without doing calculus you know this is a parabola opening downward and the maximum product will be the vertex of this parabola. did you want to do calculus on this or just find the vertex?

OpenStudy (anonymous):

calculus

OpenStudy (anonymous):

so find P' = ???

OpenStudy (anonymous):

what I m getting I don't think is right

OpenStudy (anonymous):

that's ok. ill check it. what u got? P' = ????

OpenStudy (anonymous):

p=4y

OpenStudy (anonymous):

P' = 20 - 4y this is the correct first derivative. i think you forgot to take the derivative of the first term, 20y.

OpenStudy (anonymous):

oh yeah

OpenStudy (anonymous):

now set P' = 0 and solve. what u got for y = ???

OpenStudy (anonymous):

y=-5

OpenStudy (anonymous):

i think it's positive 5. double check please.... yes/no?

OpenStudy (anonymous):

yup in is sorry

OpenStudy (anonymous):

wait no it is -5

OpenStudy (anonymous):

when you have -20=-4y when you bring the 4 over it turns to positive so the -20/4=-5

OpenStudy (anonymous):

-20=-4y is correct. so solving this equation by dividing both sides by -4, you get: \(\large \frac{-20}{-4}=\frac{-4y}{-4} \) \(\large 5=\frac{\cancel{-4}y}{\cancel{-4}} \)

OpenStudy (anonymous):

so y=5 right?

OpenStudy (anonymous):

ok yes

OpenStudy (anonymous):

ok... so y=5 and we need the x. do you know how we can get x = ???

OpenStudy (anonymous):

no

OpenStudy (anonymous):

remember x = 20 - 2y ???

OpenStudy (anonymous):

this was from your constraint equation x + 2y = 20.

OpenStudy (anonymous):

x=10

OpenStudy (anonymous):

yes... so P = xy = ???

OpenStudy (anonymous):

5(10)

OpenStudy (anonymous):

so then p=50

OpenStudy (anonymous):

ok... that's the answer BUT..

OpenStudy (anonymous):

Maximum product is50

OpenStudy (anonymous):

yes... BUT

OpenStudy (anonymous):

but what?

OpenStudy (anonymous):

how do you know that is the maximum?

OpenStudy (anonymous):

in any optimization problem you must justify that that is either a maximum or minimum...

OpenStudy (anonymous):

ok well maybe I was jumping the gun so p=5(10) so then what

OpenStudy (anonymous):

no... 50 is correct... the maximum product is 50

OpenStudy (anonymous):

ok are you sure

OpenStudy (anonymous):

you just need to justify that this product is the maximum... and yes, i'm sure...

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

how do you justify that this is the maximum?

OpenStudy (anonymous):

by checkin it right?

OpenStudy (anonymous):

that would take forever because the domain is all real numbers... maybe you can use the second derivative test? what's P'' =

OpenStudy (anonymous):

it would be um not sure I can;t get it to come out

OpenStudy (anonymous):

first derivative: P' = 20 - 4y second derivatve: P'' = -4 notice P'' is ALWAYS negative.. which means P is concave up or down?

OpenStudy (anonymous):

concave donw

OpenStudy (anonymous):

yes.... since it's concave down, the critical number is a maximum. this also verifies that in the beginning i told you this was a parabola opening downwards... you did awesome with this probem... thanks for being a great student! :)

OpenStudy (anonymous):

ok thaks